# Explaining mountain pine beetle dynamics: From life history traits to large, episodic outbreaks

### v1

@@ A.7 Integrator model
Integrator, A.7
Appendix B Additional model justification
B.1 Mass attack preference does not change
One potential shortcoming of our model is that preference only depends on tree diameter, whereas previous papers have shown that MPB’s preference for large-diameter trees increases as outbreaks progress (Boone et al.,, 2011; Howe et al.,, 2022). To investigate further, we separately fit our preference sub-model to each year of data and plotted $\beta_{1}$ (the slope of the preference $\sim$ DBH regression) against beetle pressure, as measured by the density of emerging females, and the number of emerging females per unit of susceptible tree surface area (Fig. B.1). There was no clear pattern, indicating that dynamic changes to the DBH preference are small or nonexistent. Similarly, there was no relationship between beetle pressure and the residuals of our original preference model (Fig. B.2).
[…]
Figure B.2: The mean and 95% credible intervals of the $(\text{preference})_{j,y}$ residuals — from the size-dependent preference sub-model — as a function of beetle pressure. Two metrics of beetle pressure are used: Beetles, i.e., recently emerging female beetles (units: thousands per acre), and Beetles/host, i.e., the quotient of emerging female density and within-DBH-class surface area (units: females per acre per square meter). There is no clear relationship between residual values and beetle pressure. Data from Klein et al., (1978).
There are a couple of ways to explain the absence of evidence for dynamic DBH-preferences. First, it could be that MPB’s large-tree preference “kicks in” at an early, unobserved stage of the Klein et al., outbreak. In Fig. 6. of the publication by Boone et al., (2011), the ratio of DBH of trees entered and not entered increases from 1.2 to 1.6 as an outbreak progresses from endemic to eruptive conditions. However, based on Boone et al.,’s classification scheme, the Klein et al., outbreak is approximately in the eruptive condition until the very end of the outbreak (calculations in supplementary files). Second, dynamic changes to DBH-preferences could be small. In Fig. 3D of Howe et al., (2022), the mean DBH of attacked trees only increased a few centimeters as infestation severity increased from 1 to 1000 attacked trees per hectare.
B.2 Increasing attack density does not affect population dynamics
Figure 11 in the main text shows that the density of attacks on killed trees (e.g., gallery starts per square meter) increases over time. This pattern can potentially be explained by a positive relationship between attack density and beetle population size, an assumption made by several models in the literature. To investigate further, we fit the killed-tree sub-model (Appendix A.4) to each year of data and plotted $1/\gamma$ (the model-based estimate of attack density) against beetle pressure, as measured by the density of emerging females, and the number of emerging females per unit of susceptible tree surface area (Fig. B.3). The plots are suggestive of a positive relationship if the data from 1971 are excluded, but the nature of the relationship is ultimately inconclusive.
@@ B.2 Increasing attack density does not affect population dynamics
B.3 The Allee effect depends only on beetle density
Our model states that the success probability (i.e., the probability that an attacked tree is successfully killed) only depends on beetle density. Several previous papers have modelled the success probability as an increasing function of the quotient of beetles and susceptible trees, a simplifying assumption made for mathematical tractability. In this appendix, we argue that these models 1) are unrealistic, and 2) themselves show that the quotient of beetles and trees is superfluous for modelling the Allee effect, as a practical matter.
@@ B.3 The Allee effect depends only on beetle density
Figure B.6: Estimates of $\alpha$, the proportion of susceptible trees that are attacked. The data are sourced from Boone et al., (2011, Table 1). The median of is $0.009$.
@@
Table B.1: Mean and dispersion parameters ($\mu$ and $k$ respectively) for the negative binomial distribution, fit to the distribution of attack density (units: females per square meter) across successfully and unsuccessfully attacked trees. Posterior means and 95% credible intervals endpoints are displayed. The data are sourced from Peterman, (1974); Safranyik, (1968); Shepard, (1965); Waring and Pitman, (1985). The dispersion parameter $k$ may be scale-dependent, so it is important to note the sizes of the study areas. All studies were conducted at the stand-level, but the sizes of the study areas are not directly comparable due to disparate sampling methodologies (i.e., transect vs hierarchical spatial sampling).
@@ Appendix C Model fitting details and diagnostics
To show that the marginal prior distributions had negligible influence on the posterior distribution, we computed the posterior contraction.
@@
a parameter-specific relative influence of the data on parameter estimation. For nearly every parameter in every model, the prior contraction was above 0.9, implying negligible prior influence. The exception is the Allee threshold parameter ($a$) in the complex outbreak dynamics threshold. The posterior contraction analysis is provided in the R markdown file model_diagnostics.Rmd and its output model_post_contract_printoff.txt, in the supplementary files.
@@
$\sigma_{\text{pref0}}$ pref. baseline noise 0.630 0.0822 0.487 0.810
$\sigma_{\text{pref1}}$ pref. sampling noise 0.152 0.113 0.00478 0.412
@@
$\sigma_{\text{dead}}$ dead tree noise 0.00508 0.000668 0.00399 0.00660
@@
$\sigma_{\text{exit}}$ emerge noise 3.61 1.54 1.79 7.47
$s_{0}$ max dispersal survival 0.474 0.0676 0.363 0.627
$\sigma_{0}$ residual noise 365 89.7 224 571
$\lambda_{\text{resid}}$ dispersal threshold 0.652 0.0339 0.582 0.705
@@
$\sigma_{\text{tree}}$ tree abund. noise 0.482 0.211 0.240 1.02
@@
$-\beta_{0}/\beta_{1}$ susc. DBH threshold 7.43 0.643 6.04 8.43
$1/\gamma$ attack density 66.2 3.78 59.2 74.1
$\zeta_{0}+\zeta_{1}*10$ emerge density (10 in. DBH) 244 25.4 192 297
Table C.1: Summary statistics for all model parameters, including the Mean, standard deviation, and parameter credible intervals ($\text{CI}_{2.5\%}$, $\text{CI}_{97.5\%}$). Here, the parameters come from the full hierarchical model (Appendix A) fit to the Klein et al., dataset. The table also includes derived quantities, including the DBH of the smallest susceptible trees, the model-estimated attack density (gallery starts per square meter), and the emergence density (females per square meter) on a tree with a 10-inch DBH, which is approximately the average DBH of attacked trees in the Klein et al., dataset. Finally, the table contains fixed parameters, whose values were obtained via literature search. For parameter units, and the summary statistics of sub-model noise parameters, see supplementary Table A.1.
@@
a Allee threshold 2492.2411175 569 1631.332 3809.797
$\sigma_{\text{Allee}}$ Pr. success noise 0.207 0.0300 0.158 0.275
@@
Appendix D Supplementary Figures
Figure D.1: Approximate dispersal mortality across time in the Klein et al., data. The dispersal mortality was estimated as the quotient of gallery starts and preceding emergence holes. The approximation will be better when emergence density is larger, because the Allee effect will be less important, and local population dynamics — compared to dispersal — will be more important.
Figure D.2: Population maps provide evidence for compensatory density dependence at the tree level. The data comes from Alberta’s disk surveys (described by author N/A,, 2020, and provided by Goodsman et al.,, 2018). The estimate for attack density at time $t+1$ involves the constants $(2/3)$ (the average proportion of the brood that will be female) and $0.3$ (an estimate of dispersal-phase mortality from the Klein et al., (1978) dataset, see Fig. D.1).

### v2

@@ B.2 Mass attack preference does not change
Figure B.2: The mean and 95% credible intervals of the $(\text{preference})_{j,y}$ residuals — from the tree size-dependent preference sub-model — as a function of beetle pressure. Two metrics of beetle pressure are used: Beetles, i.e., recently emerging female beetles (units: thousands per acre), and Beetles/host, i.e., the quotient of emerging female density and within-DBH-class surface area (units: females per acre per square meter). There is no clear relationship between residual values and beetle pressure. Data from Klein et al., (1978).
There are a couple of ways to explain the absence of evidence for dynamic DBH-preferences. First, it could be that MPB’s large-tree preference “kicks in” at an early, unobserved stage of the Klein et al., outbreak. In Fig. 6. of the publication by Boone et al., (2011), the ratio of DBH of trees entered and not entered increases from 1.2 to 1.6 as an outbreak progresses from endemic to eruptive conditions. However, based on Boone et al.,’s classification scheme, the Klein et al., outbreak is approximately in the eruptive condition until the very end of the outbreak (calculations in supplementary files). Second, dynamic changes to DBH-preferences could be small. In Fig. 3D of Howe et al., 2022b , the mean DBH of attacked trees only increased a few centimeters as infestation severity increased from 1 to 1000 attacked trees per hectare.
B.3 Increasing attack density does not affect population dynamics
Figure A.9 in the main text shows that the density of attacks on killed trees (e.g., gallery starts per square meter) increases over time. This pattern can potentially be explained by a positive relationship between attack density and beetle population size, an assumption made by several models in the literature. To investigate further, we fit the killed-tree sub-model (Appendix A.6) to each year of data and plotted $1/\gamma$ (the model-based estimate of attack density) against beetle pressure, as measured by the density of emerging females, and the number of emerging females per unit of susceptible tree surface area (Fig. B.3). The plots are suggestive of a positive relationship if the data from 1971 are excluded, but the nature of the relationship is ultimately inconclusive.
@@ B.3 Increasing attack density does not affect population dynamics
B.4 Leave one year out cross validation
The statistical model successfully predicts in-sample data (Fig. B.5, panel A) and out-of-sample data. The exception is that the out-of-sample predictive intervals for 1970 are wide (Fig. B.5, panel B), which upon further investigation, was found to be the result of uncertain estimates of the stand-dependent dispersal mortality parameters, $\lambda_{0}$ and $\lambda_{1}$ (Fig. B.5, panel C).
Figure B.5: Observations of successful attack density in the Klein et al., dataset (operationalized as gallery starts), with superimposed model-based predictive intervals (PIs). Panel A shows in-sample predictions for the full statistical model; panel B shows out-of-sample, leave-one-year-out (LOYO) predictions, where the model is not trained on the focal year’s data; and panel C shows LOYO predictions when the stand-dependent dispersal mortality parameters ($\lambda_{0}$ and $\lambda_{1}$) are fixed at the medians of their respective marginal posterior distributions (see Table 1). Panels B & C demonstrate that data from a single beetle generation (1969–1970) has an overwhelming influence on the estimation of $\lambda_{0}$ and $\lambda_{1}$, but that all other parameter estimates are robust.
The estimation of stand-dependent dispersal mortality parameters is highly dependent on the data from 1969–1970. This would be indicative of overfitting (i.e., fitting the noise) if the decrease in attack density in 1970 was the result of noise, rather than stand-dependent dispersal mortality. The obvious source of statistical noise is the weather, but the weather leading up to the summer of 1970 was unexceptional (Figs. B.7 & B.6). Taken together, evidence for the presence of stand-dependent dispersal mortality (Section 1.2.2) and the absence of anomalous weather patterns imply that the model did not overfit.
Figure B.6: In 1970, emergence density was high but subsequent attack density was low. Anomalous weather conditions during the flight period could be responsible for this pattern, but the relevant weather variables were unexceptional in 1970 (the bold, teal line). The maximum summer temperature was consistently greater than 18.3°C, the lower limit for beetle flight (McCambridge,, 1971). The estimated wind speed often exceeded 8 km/hour, the proposed upper threshold for beetle flight (Gray et al.,, 1972); however, the wind speeds in 1970 were not anomalous in comparison to the other outbreak years. Weather data was obtained from the Biosim program (Régnière et al.,, 2014).
Figure B.7: In 1970, emergence density was high but subsequent attack density was low. The decline could theoretically be attributed to poor environmental conditions that weakened the beetles, despite not being lethal (because they still emerged). However, the developmental period leading up to 1970 had unexceptional weather conditions. The minimum winter temperature is related to larval mortality (here, sub-lethal damage), and degree days are related to emergence and univoltinism. The minimum winter temperature is approximately -30°C, which is substantially greater than the proposed threshold of -40°C (Safranyik,, 1998) for significant overwintering mortality. There are approximately 1250 degree days from Aug. 1 to Jul. 31, far beyond the proposed threshold of 833 for the population to be univoltine (Reid,, 1962). Weather data was obtained from the Biosim program (Régnière et al.,, 2014).
B.5 The Allee effect depends only on beetle density
Our model states that the success probability (i.e., the probability that an attacked tree is successfully killed) only depends on beetle density. Several previous papers have modeled the success probability as an increasing function of the quotient of beetles and susceptible trees, a simplifying assumption made for mathematical tractability. In this appendix, we argue that these models 1) are unrealistic, and 2) themselves show that the quotient of beetles and trees is superfluous for modeling the Allee effect, as a practical matter.
@@ B.5 The Allee effect depends only on beetle density
Figure B.9: Estimates of $\alpha$, the proportion of susceptible trees that are attacked. The data are sourced from Boone et al., (2011, Table 1). The median is $0.009$.
@@
Table B.1: Mean and dispersion parameters ($\mu$ and $k$ respectively) for the negative binomial distribution, fit to the distribution of attack density (units: females per square meter) across successfully and unsuccessfully attacked trees. Posterior means and 95% credible intervals endpoints are displayed. The data are sourced from Peterman, (1974); Safranyik, (1968); Shepard, (1965); Waring and Pitman, (1985). The dispersion parameter $k$ may be scale-dependent, so it is important to note the sizes of the study areas. All studies were conducted at the stand level, but the sizes of the study areas are not directly comparable due to disparate sampling methodologies (i.e., transect vs hierarchical spatial sampling).
@@ Appendix C Model fitting details and diagnostics
To show that the marginal prior distributions had negligible influence on the posterior distribution, we computed the posterior contraction,
@@
This is a parameter-specific metric that shows the influence of the data on parameter estimation. For nearly every parameter in every model, the prior contraction was above 0.9, implying negligible prior influence. The exception is the Allee threshold parameter in the complex outbreak dynamics threshold. The posterior contraction analysis is provided in the R markdown file model_diagnostics.Rmd and its output model_post_contract_printoff.txt, in the supplementary files.
The low value of the posterior contraction for the parameter $a$ indicates that the Klein et al., data is not informative with respect to estimating the Allee effect. This is unsurprising, as discerning the Allee effect is challenging without data from trees that were attacked but not successfully killed (see Section 1.2.1), data which is absent in Klein et al.,’s 1978 publication. However, the Allee threshold estimate is not devoid of information, since the prior distribution was derived from the Allee effect sub-model fit to the Boone et al., dataset (see Table C.2 & Appendix A.8).
@@
$\sigma_{\text{pref0}}$ pref. baseline noise 0.628 0.0830 0.484 0.814
$\sigma_{\text{pref1}}$ pref. sampling noise 0.154 0.112 0.00663 0.417
@@
$\sigma_{\text{dead}}$ dead tree noise 0.00508 0.000657 0.00398 0.00653
@@
$\sigma_{\text{exit}}$ emerge noise 3.61 1.52 1.76 7.42
$s_{0}$ max dispersal survival 0.477 0.0701 0.363 0.639
$\sigma_{0}$ residual noise 363 88.7 225 568
$\lambda_{\text{resid}}$ dispersal threshold 0.652 0.0353 0.579 0.708
@@
$-\beta_{1}/\beta_{0}$ susc. DBH threshold 7.46 0.666 5.98 8.47
$1/\gamma$ attack density 66.3 3.88 59.4 74.4
$\zeta_{0}+\zeta_{1}*10$ emerge density (10 in. DBH) 244 25.1 191 296
$K-\sum_{t=0}^{17}(1-m)^{t}mK$ Overstory density 338 86.0 193 529
Table C.1: Summary statistics for all model parameters, including the Mean, standard deviation, and parameter credible intervals ($\text{CI}_{2.5\%}$, $\text{CI}_{97.5\%}$). Here, the parameters come from the full hierarchical model (Appendix A) fit to the Klein et al., dataset. The table also includes derived quantities, including the DBH of the smallest susceptible trees, the model-estimated attack density (gallery starts per square meter), and the emergence density (females per square meter) on a tree with a 10-inch DBH, which is approximately the average DBH of attacked trees in the Klein et al., dataset. Finally, the table contains fixed parameters, whose values were obtained via literature search. For parameter units, and the summary statistics of sub-model noise parameters, see supplementary Table LABEL:tab:model_notation.
@@
a Allee threshold 2492 569 1631 3809
$\sigma_{\text{Allee}}$ Pr. success noise 0.207 0.03 0.158 0.275
@@
Appendix D Landscape analysis of bark beetle damage
Here, we analyze the relationship between bark beetle damage and host tree biomass across western North America, resulting in Figure 9 (in the main text). We focused on entirety of British Columbia and regions 1–6 of the United States (defined by the United States Forest Service), as these areas contain the majority of bark beetle data and initiated data collection earlier than other regions. The temporal scope of our analysis begins with the earliest start date for Aerial Detection Survey (ADS) data in each region, or 1962 when ADS data preceded this date, since surveys were not comprehensive before 1962. Specifically, the start years are: Region 1 (1962), Region 2 (1994), Region 3 (1997), Region 4 (1991), Region 5 (1996), Region 6 (1962), British Columbia (1962), and Alberta (1998). All regions have an end date of 2021.
Bark beetle damage is estimated using publicly available Aerial Detection Survey (ADS) data. This comprehensive dataset is collected annually through aerial surveillance of all coniferous forests within defined regions. Trained surveyors fly over these areas, delineating polygons that represent the extent of observed forest damage while simultaneously estimating damage severity (e.g., percentage of trees affected within each polygon). Due to variations in severity metrics across space and time, we simplified our analysis by measuring bark beetle damage as the total affected area, summing affected polygon areas across all years and regions. Host tree species information primarily comes from Wood, (1982), with additional references for the North American spruce beetle (USDA Forest Service,, 2011; British Columbia Government, nd, ) and the southern pine beetle (North Carolina Forest Service, nd, ; USDA Forest Service,, 2009). Host-trees are given in Table LABEL:tab:host_trees. For host-tree biomass data, we utilized multiple sources. In British Columbia, we used species-specific estimates of total live above-ground biomass from Beaudoin et al., (2014). In the United States, we calculated biomass by multiplying relative abundances from Individual Tree Species Parameter (ITSP) maps (Ellenwood et al.,, 2015) by total biomass data from Blackard et al., (2008).
Our methodology involved classifying all bark beetles in the ADS data into three categories: irruptive (also referred to as excitable or aggressive), pulse-driven (semi-aggressive or opportunistically aggressive), and non-outbreaking (non-aggressive or lower-stem beetles). This classification is based on subjective judgments gleaned from the literature, as there is no single diagnostic feature of irruptive bark beetle species. We designated the following species as irruptive: mountain pine beetle, European spruce beetle, southern pine beetle, western pine beetle, and western balsam bark beetle, supported by Singh et al., (2024), Howe et al., 2022a , Howe et al., 2022b , Lantschner and Corley, (2023), Biedermann et al., (2019), Koontz et al., (2021), and Weed et al., (2015). Pulse-driven species were defined as species that could outbreak but were not known to be irruptive (using Table 1 in Lantschner and Corley, (2023), Table 8.1 in Six and Bracewell, (2015), and Table 4.1 in Weed et al., (2015)) Non-outbreaking species are provided by Howe et al., 2022b ; all remaining beetle species in the ADS data were assumed to be non-outbreaking.
BB Common Name BB Genus Species Host Common Name Host Genus Species Notes
roundheaded pine beetle Dendroctonus adjunctus Mexican white pine Pinus ayacahuite Mexico & CentAm
Mexican mountain pine Pinus hartwegii Mexico & CentAm
Chihuahua pine Pinus leiophylla Mexico & S. NM/AZ
Montezuma pine Pinus montezumae Mexico & CentAm
ponderosa pine Pinus ponderosa BC & W. USA
smooth-bark Mexican pine Pinus pseudostrobus Mexico & CentAm
thinleaf pine Pinus maximinoi Mexico & CentAm
western pine beetle Dendroctonus brevicomis ponderosa pine Pinus ponderosa BC & W. USA
Jeffrey pine beetle Dendroctonus jeffreyi Jeffrey pine Pinus jeffreyi CA, NV, OR, NW Mexico
ponderosa pine Pinus ponderosa (very rare) BC & W. USA
lodgepole pine beetle Dendroctonus murrayanae jack pine Pinus banksiana CAN: BC-NS, USA: MN-ME
lodgepole pine Pinus contorta W. NorAm
eastern white pine Pinus strobus NL to great lakes
mountain pine beetle Dendroctonus ponderosae Engelmann spruce Picea engelmanii W. NorAm
whitebark pine Pinus albicaulis W. Can & W. USA
foxtail pine Pinus balfouriana CA only
jack pine Pinus banksiana
lodgepole pine Pinus contorta
Coulter/big-cone pine Pinus coulteri S. CA & NW Mexico
pinyon pine Pinus edulis SW USA
limber pine Pinus flexilis W NorAm
Jeffrey pine Pinus jeffreyi (rare)
sugar pine Pinus lambertiana OR to NW Mexico
single-leaf pinyon pine Pinus monophylla ID to NW Mexico
western white pine Pinus monticola W. Can & USA
ponderosa pine Pinus ponderosa BC & W. USA
Chihuahua white pine Pinus strobiformis Mexico & SW USA
douglas-fir beetle Dendroctonus pseudotsugae western larch/tamarack Larix occidentalis (rare) NW Can & USA
bigcone spruce/bigcone Douglas-fir Pseudotsuga macrocarpa CA only
Douglas-fir Pseudotsuga menziesii W. NorAm
western hemlock Tsuga heterophylla NW Can and USA incl AK
spruce beetle Dendroctonus rufipennis spruce species All Picea species in its range (mostly engelmann and white spruce)
eastern larch beetle Dendroctonus simplex eastern larch/tamarack Larix laricina All Can, AK & NE USA
red turpentine beetle Dendroctonus valens white fir Abies concolor W. USA
eastern larch/tamarack Larix larciana
Norway spruce Picea excelsa Introduced NE USA & Can
white spruce Picea glauca All Can, AK, & N. USA
red spruce Picea rubens NE USA, Quebec-NS
lodgepole pine Pinus contorta
Coulter/big-cone pine Pinus coulteri
shortleaf pine Pinus echinata SE USA
pinyon pine Pinus edulis
Mexican mountain pine Pinus hartwegii
Jeffrey pine Pinus jeffreyi
sugar pine Pinus lambertiana
Lawson’s pine Pinus lawsonii Mexico
Chihuahua pine Pinus leiophylla
western white pine Pinus monticola
Sierra lodgepole pine Pinus murrayana
Mexican yellow pine Pinus oocarpa Mexico & CentAm
ponderosa pine Pinus ponderosa BC & W. USA
smooth-bark Mexican pine Pinus pseudostrobus
Monterey pine Pinus radiata Coastal CA & Mexico
red pine Pinus resinosa E Can & NE USA
pitch pine Pinus rigida E. USA
California foothill pine Pinus sabiniana California only
Chihuahua white pine Pinus strobiformis
eastern white pine Pinus strobus
Scotch pine Pinus sylvestris Introduced
thinleaf pine Pinus maximinoi
Virginia pine Pinus virginiana E. USA
NA Dryocoetes autographus fir species Abies spp.
Engelmann spruce Picea engelmanii
white spruce Picea glauca
lodgepole pine Pinus contorta
western white pine Pinus monticola
eastern white pine Pinus strobus
Douglas-fir Pseudotsuga menziesii
western hemlock Tsuga heterophylla
western balsam bark beetle Dryocoetes confusus pacific silver fir Abies amabilis (rare) W. BC, WA, OR
white fir Abies concolor (rare)
subalpine fir Abies lasiocarpa
Engelmann spruce Picea engelmanii
pinyon ips Ips confusus Colerado pinyon pine Pinus edulis
single-leaf pinyon pine Pinus monophylla
emarginate Ips Ips emarginatus lodgepole pine Pinus contorta (rare)
Jeffrey pine Pinus jeffreyi
western white pine Pinus monticola
ponderosa pine Pinus ponderosa BC & W. USA
blue spruce engraver Ips hunteri Engelmann spruce Picea engelmanii (rare)
blue spruce Picea pungens
California fivespined Ips Ips paraconfusus knobcone pine Pinus attenuata
lodgepole pine Pinus contorta (rare)
Coulter/big-cone pine Pinus coulteri
Jeffrey pine Pinus jeffreyi
sugar pine Pinus lambertiana
ponderosa pine Pinus ponderosa BC & W. USA
northern spruce engraver Ips perturbatus white spruce Picea glauca
pine engraver Ips pini jack pine Pinus banksiana
lodgepole pine Pinus contorta
Jeffrey pine Pinus jeffreyi
ponderosa pine Pinus ponderosa BC & W. USA
eastern white pine Pinus strobus
western cedar bark beetle Phloeosinus spp. Rocky Mountain juniper Juniperus scopulorum
Utah juniper Juniperus osteosperma
one-seed juniper Juniperus monosperma
NA Pityophtorus pseudotsugae white fir Abies concolor
grand fir Abies grandis
subalpine fir Abies lasiocarpa
red fir Abies magnifica
Engelmann spruce Picea engelmanii
sugar pine Pinus lambertiana
Douglas-fir Pseudotsuga menziesii
western hemlock Tsuga heterophylla
fir root bark beetle Pseudohylesinus granulatus pacific silver fir Abies amabilis
grand fir Abies grandis
subalpine fir Abies lasiocarpa
red fir Abies magnifica
noble fir Abies procera
western hemlock Tsuga heterophylla
Douglas-fir pole beetle Pseudohylesinus nebulosus Douglas-fir Pseudotsuga menziesii
western hemlock Tsuga heterophylla (rare)
silver fir beetle Pseudohylesinus sericeus pacific silver fir Abies amabilis
grand fir Abies grandis
noble fir Abies procera
Douglas-fir Pseudotsuga menziesii
western hemlock Tsuga heterophylla
Douglas-fir engraver Scolytus unispinosus Douglas-fir Pseudotsuga menziesii
fir engraver Scolytus ventralis white fir Abies concolor
grand fir Abies grandis
red fir Abies magnifica
Mediterranean oak borer Xyleborus monographus blue oak Quercus douglasii
valley oak Quercus lobata
southern pine beetle Dendroctonus frontalis shortleaf pine Pinus echinata
loblolly pine Pinus taeda
pitch pine Pinus rigida
pond pine Pinus serotina
Virginia pine Pinus virginiana
Appendix E Additional supplementary figures